OpenStudy (anonymous):

find the equation of the tangent line to the graph of y = x + tanx at the point (pi , pi). write in slope intercept form

OpenStudy (anonymous):

i used the first derivative test and i got 1+sec^2(x)

OpenStudy (anonymous):

idk what to do next

myininaya (myininaya):

find the slope by pluggin in pi

myininaya (myininaya):

into your derivative

OpenStudy (anonymous):

so in the calculator it would look 1+ 1/(cos(pi)^2 right? i got 2 for the slope

myininaya (myininaya):

cos(pi)=-1 cos^2(pi)=1 so 1/cos^2(pi)=1/1=1 so yeah 1+1 is 2

myininaya (myininaya):

the slope if 2

myininaya (myininaya):

y-y1=m(x-x1)

myininaya (myininaya):

you know m=2

myininaya (myininaya):

and you also can find a point on the curve and that is on the tangent line

OpenStudy (anonymous):

yeah i got y-pi= 2(x-pi)

myininaya (myininaya):

oh yeah it was already given was thinking you had to plug in pi just to figure out it was pi as output too :)

OpenStudy (anonymous):

:D thanks

myininaya (myininaya):

now put that in whatever form you want and can but that is the tangent line in point slope form

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so for example if in the quiz they ask me the same thing but for a different question like y= x+cotx

OpenStudy (anonymous):

i just use the first derivative

OpenStudy (anonymous):

and evaluate for pi right?

OpenStudy (anonymous):

if (pi,pi) was given like this time

myininaya (myininaya):

well if it says to find the tangent line at x=pi yeah

myininaya (myininaya):

but cot(pi)=cos(pi)/sin(pi) which doesn't exist

myininaya (myininaya):

so that question actually wouldn't make sense

OpenStudy (anonymous):

hehe :D

OpenStudy (anonymous):

you know what i meant though. thanks alot

OpenStudy (anonymous):

finish what you are typing i got a last question :)

myininaya (myininaya):

if we are asked to find the tagent line to y=x+cot(x) at (a,b) then first step find slope=1-csc^2(x) find slope (@x=a)=1-csc^2(a) <--this is slope so we have at x=a, y=a+cot(a)=b so the tangent line here would be y-b=m(x-a) y-a-cot(a)=(1-csc^2(a))(x-a) y=(1-csc^2(a))(x-a)+a+cot(a) y=(1-csc^2(a))x-a(1-csc^2(a)+a+cot(a) y=(1-csc^2(a))x-a-acsc^2(a)+a+cot(a) y=(1-csc^2(a))x-acsc^2(a)+cot(a)

OpenStudy (anonymous):

there is a question in my practice quiz like this y=sin(sin(sinx)

myininaya (myininaya):

find y'?

OpenStudy (anonymous):

no differentiate

myininaya (myininaya):

?

myininaya (myininaya):

don't find y'?

myininaya (myininaya):

or do?

OpenStudy (anonymous):

it says differentiate that function

myininaya (myininaya):

so that means we need to find y'

myininaya (myininaya):

or to find the derivative

OpenStudy (anonymous):

yep i wanted to know if there was a difference between that function and sinx.sinx.sinx

myininaya (myininaya):

\[y=f(u(x)) \\ y'=f'(z)|_{z=u(x)} \cdot u'(x)\] do you know the chain rule?

OpenStudy (anonymous):

yes

myininaya (myininaya):

yes sin(x)*sin(x)*sin(x) is a product

myininaya (myininaya):

sin(sin(sin(x))) is a function inside a function inside a function

OpenStudy (anonymous):

hmm

myininaya (myininaya):

one requires product rule the other requires chain rule

myininaya (myininaya):

they are not the same function at all

OpenStudy (anonymous):

so for the product one would you get this

OpenStudy (anonymous):

cosxsinxsinx+ sinxcosxsinx+sinxsinxcosx?

myininaya (myininaya):

for example at x=pi/2 sin(pi/2)*sin(pi/2)*sin(pi/2) =1*1*1=1 but sin(sin(sin(pi/2))) =sin(sin(1)) approx sin(.84147) approx .745623 but anyways... and yes to that question about the product or you could write it as sin^3(x) and then differentiate either way \[\frac{d}{dx}\sin^3(x)=3\sin^2(x)\cos(x) \] that is exactly what you get doing it the way you chose

OpenStudy (anonymous):

i see thanks :D

myininaya (myininaya):

but the chain rule one can you do that one?

OpenStudy (anonymous):

would that be cos(sin(sinx).cosx(sinx).cosx

myininaya (myininaya):

\[y=f(u(x)) \\ y'=f'(z)|_{z=u(x)} \cdot u'(x) \\ y=\sin(\sin(\sin(x)))\] let f(z)=sin(z) and u(x)=sin(sin(x)) so f'(z)=cos(z) so f'(u(x))=cos(sin(sin(x)) and we could use the chain rule again for the sin(sin(x)) part so u(x)=g(h(x)) u'(x)=g'(h(x))*h'(x) u'(x)=cos(sin(x))*cos(x) so yeah you get y'=cos(sin(sin(x))*cos(sin(x))*cos(x)

myininaya (myininaya):

good job

myininaya (myininaya):

i know that is probably longer of an explanation then you wanted just trying to be thorough

OpenStudy (anonymous):

that's perfect thanks